OpenStudy (anonymous):

Change 2x + 5y = –30 to slope-intercept form. You must show all work

OpenStudy (anonymous):

Do you know what slop-intercept form looks like?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Slope-intercept is y=mx+b

OpenStudy (anonymous):

okay and how does this work

OpenStudy (anonymous):

First try rearranging the equation to fit this form

OpenStudy (anonymous):

im so sorry but i dont know how i just got transferred in to this school last week and they where already teaching this when i got their

OpenStudy (anonymous):

@Karebear_Karl you their

OpenStudy (anonymous):

Okay, y=mx+b

OpenStudy (anonymous):

Re-arrange it.

OpenStudy (anonymous):

but how do i rearrange it

OpenStudy (anonymous):

Just move each number around.

OpenStudy (anonymous):

The coefficiant of each variable stays on.

OpenStudy (anonymous):

This is what it should look like 5y=2x-30. You need to get y by it's self... how do you think you could do that?

OpenStudy (anonymous):

divide by 5 to both sides

OpenStudy (anonymous):

And what do you get?

OpenStudy (anonymous):

y=2x-30/5

OpenStudy (anonymous):

Are you writing or typing the answer?

OpenStudy (anonymous):

When you type it, type it with perentheses, like y=(2x-30)/5

OpenStudy (anonymous):

Now what is 30 divided by 5...

OpenStudy (anonymous):

6

OpenStudy (anonymous):

This should be your answer... y = -2/5x - 6. Sorry I'm not the best at explaining math..

OpenStudy (anonymous):

but how do i put it in slope intercept form

OpenStudy (anonymous):

slope is -2/5 and y intercept is (0,-6) That is y intercept form. Y intercept form equation is y=mx+b.

OpenStudy (anonymous):

Our first step was changing it to y intercept form. :)

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

You are very welcome. Good luck! :)